OpenStudy (mateeldaa):

cos theta=7/25, sin theta<0 find sin 2theta

OpenStudy (anonymous):

\[ \cos^2(x) + \sin^2(x) = 1\\ \sin( 2 a) = 2 \sin(a) \cos(a) \]

OpenStudy (mateeldaa):

that is right, but I don't understand what sin (a) is...

OpenStudy (anonymous):

\[ \cos^2(\theta) + \sin^2(\theta) = 1\\ \sin^2(\theta) = 1-\cos^2(\theta)\\ \sin^2(\theta) = 1-\left( \frac 7 {25}\right)^2= 1- \frac {49}{625}=\frac{576}{625} =\left(\frac{24}{25}\right)^2\\ \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (mateeldaa):

2*(24/25)*(7/25)=.5376 sin theta<0= -.5376 is this right?

OpenStudy (anonymous):

No \[ \sin(\theta) =-\frac{24}{25} \]

OpenStudy (anonymous):

\[ \sin^2(\theta) = \left(\frac{24}{25}\right)^2\\ \sin(\theta) = \pm \frac {24}{25} \]

OpenStudy (anonymous):

We take the negative side since it was given to be <0

OpenStudy (anonymous):

Got it? Can you continue the problem?

OpenStudy (anonymous):

\[ \sin(2 \theta) = 2 \sin(\theta) cos(\theta) \]

OpenStudy (mateeldaa):

Yes, I got it. So then 2*(-24/25)*(7/25)= -336/625

OpenStudy (anonymous):

Excellent

OpenStudy (mateeldaa):

Thank you so much for your help! :)